OpenStudy (anonymous):

\[\frac{x^2+(x+1)(x+2)}{5x+1}=6\]

OpenStudy (anonymous):

help solve the equation

OpenStudy (anonymous):

\[\frac{x^2-(x+1)(x+2)}{5x+1}=6\]

OpenStudy (anonymous):

\[\frac{x^2-(x+1)(x+2)}{5x+1}=6\] this is the correct equation

OpenStudy (anonymous):

I would first multiply both sides by 5x + 1 x^2 - (x+1)(x+2) = 30x + 6 Then expand x^2 - x^2 - 3x - 2 = 30x + 6 *note the "-" signs as it has to distribute to all the terms

OpenStudy (anonymous):

simplifying further: -3x - 2 = 30x + 6 -8 = 33x -8/33 = x

OpenStudy (anonymous):

-3x-2=30x+6 -3x-30x=6+2 -33x=8 x=8/(-33) thanks i got it